OpenStudy (anonymous):

A metal washer 1-inch in diameter is pierced by a 1/2 - inch hole. What is the volume of the washer if it is 1/8 - inch thick?

OpenStudy (anonymous):

V=A*T T=1/8 A=PI(1-(1/2)^2) V=(1/8)*(3/4)*PI

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

no problem.